OpenStudy (anonymous):

how to get a standard deviation from a set of numbers, simple explanation please.

OpenStudy (amistre64):

hmmm, its variance squared

OpenStudy (anonymous):

ya, square root of variance

OpenStudy (amistre64):

find the mean subtract the mean from all the points and square the results add up all the results ... and divide by how many there are if its a population, or by one less if its just a sample

OpenStudy (amistre64):

then sqrt that result i think

OpenStudy (amistre64):

var=\[\frac{\sum(x-\bar x)^2}{N(or\ n-1)}\] right?

OpenStudy (amistre64):

sqrt that to get sd

OpenStudy (anonymous):

i don't think its over n

OpenStudy (amistre64):

i wouldnt know, i cant see if you got population or sample

OpenStudy (anonymous):

and stand dev is the square root of variance

OpenStudy (amistre64):

the sample has to do with the degrees of freedom i think

OpenStudy (amistre64):

thats the one, sd = sqrt(var)

OpenStudy (anonymous):

so, add up all the differences for each value from the mean and square the sum

OpenStudy (anonymous):

then....

OpenStudy (amistre64):

add up all the squares of the differences from the mean

OpenStudy (amistre64):

divide it by the correct n or n-1 depending on technical stuff

OpenStudy (amistre64):

then sqrt it

OpenStudy (anonymous):

ahhh, ok

OpenStudy (amistre64):

the sqrt at the end is spose to counteract the square at the start.

OpenStudy (anonymous):

ok, i can do calc all day, can't stand stats

OpenStudy (amistre64):

stats was fun last semester